# Continuous K-space Recovery Network with Image Guidance for Fast MRI Reconstruction

### v1

@@ IV-C2 Efficacy of recovering k-space
To further analyze the effectiveness of our approach in the recovery of k-space, we present the results of k-space reconstruction, as shown in Fig. 7. Specifically, we show the sample of reconstructed images and absolute differences (10×10$\times$10 ×) of standardized pixel intensities between reconstructed k-space and GT ones. It can be observed that our method achieves a generally low error in the recovered spectrum. Corresponding to the MRI images, it is evident that our approach excels in recovering high-frequency details and edges, while other methods tend to exhibit issues of over-smoothing. This further substantiates the importance of k-space recovery for the precise reconstruction of MRI images and the effectiveness of our proposed method.
V CONCLUSION

### v2

@@ 4.3.2 Efficacy of Recovering K-space
To further analyze the effectiveness of our approach in the recovery of k-space, we present the results of k-space reconstruction, as shown in Figure. 8. Specifically, we show the sample of reconstructed images and absolute differences (10×10$\times$10 ×) of standardized pixel intensities between reconstructed k-space and GT ones. It can be observed that our method achieves a generally low error in the recovered spectrum. Corresponding to the MRI images, it is evident that our approach excels in recovering high-frequency details and edges, while other methods tend to exhibit issues of over-smoothing. This further substantiates the importance of k-space recovery for the precise reconstruction of MRI images and the effectiveness of our proposed method.
Figure 8: Visual comparison results with k-space error maps.
Table 7: Efficiency comparison on CC359 dataset. (1D Cartesian 20%)
Method FLOPs Param. PSNR↑ SSIM↑ NMSE↓ LPIPS↓
SwinMR 398G 11.40M 31.11 0.9006 0.0124 0.0413
ReconFormer 342G 1.14M 32.79 0.9189 0.0072 0.0373
Ours 138G 13.89M 33.06 0.9234 0.0069 0.0362
4.3.3 Analysis of Training Efficiency
The results of the training efficiency comparison are reported in Table 7. SwinMR obtained a small number of parameters but with a large computational complexity (FLOPs=$398$G) and suboptimal reconstruction results. ReconFormer used a local pyramidal but global columnar transformer structure, which achieved a lightweight parameter of $1.14$M via its recurrent nature. However, the larger computational complexity (FLOPs = $342$G) hinders its training and testing process. Our IGKR-Net outperformed other methods and obtained an optimal balance between the number of parameters and computational complexity.
5 Discussion and Future Work
This paper introduces IGKR-Net, a network that addresses the challenging problem of k-space recovery caused by under-sampling, using INR as a new perspective. Furthermore, by introducing an image guidance module that extracts information from low-quality MRI images and proposing a tri-attention refinement module that enhances the resulting image quality, IGKR-Net ultimately achieves accurate MRI reconstruction. Experimental results demonstrate our method’s effectiveness, showing superior performance compared to existing MRI reconstruction techniques on various datasets. In future work, we will also extend IGKR-Net to other recovery tasks, such as cardiac MRI.
6 Conclusion